# Connecting the Dots: Is Mode-Connectedness the Key to Feasible Sample-Based Inference in Bayesian Neural Networks?

### v1

@@ Abstract
A major challenge in sample-based inference (SBI) for Bayesian neural networks is the size and structure of the networks’ parameter space. Our work shows that successful SBI is possible by embracing the characteristic relationship between weight and function space, uncovering a systematic link between overparameterization and the difficulty of the sampling problem. Through extensive experiments, we establish practical guidelines for sampling and convergence diagnosis. As a result, we present a Bayesian deep ensemble approach as an effective solution with competitive performance and uncertainty quantification.
@@ 1 Introduction
Bayesian neural networks (BNNs) represent a principled solution to the problem of probabilistic deep learning. In the absence of analytically tractable solutions, Bayesians traditionally rely on sample-based inference (SBI) as it can in theory recover the true posterior—up to a Monte Carlo error—and requires no (restrictive) assumptions on the posterior distributional family. Consequently, SBI has enormous potential for uncertainty quantification in BNNs (Farquhar et al., 2020; Izmailov et al., 2021; Wiese et al., 2023). However, recent research has focused on local approximations that may not fully capture the multimodality of BNN posteriors, as noted in Alexos et al. (2022); Arbel et al. (2023). Such approaches potentially overlook significant portions of the posterior density, which would limit the comprehensiveness of uncertainty quantification. The main reason behind the reluctance to adopt SBI appears to be its perceived computational demands, as discussed in Papamarkou et al. (2022); Sharma et al. (2023b). Indeed, Izmailov et al. (2021) show that SBI is possible for large state-of-the-art (SOTA) networks but comes at a high cost. Yet, scalable software solutions (e.g., JAX, Bradbury et al., 2018) and methodological adaptations (Nemeth & Fearnhead, 2021) have made SBI increasingly accessible.
While technological progress continues unabated, work studying the peculiarities of SBI is limited. One reason for this research gap, and possibly the most fundamental problem for BNNs, is their large number of parameters. The resulting challenges are uncharted territory for established Bayesian workflow routines such as the sampling strategy, the choice of prior, and convergence monitoring. It is tempting to conclude that parameter space inference should simply be discarded in favor of addressing uncertainties directly in the function space (see, e.g., Tran et al., 2022), but this perspective risks suspending valuable insights.
@@
Our Contributions. Among the discussions about the merits of parameter space inference (see also Appendix A), our paper shows that SBI can be successfully leveraged for BNNs. We argue that accounting for the idiosyncratic relationship between weight and function space is key to affordable SBI. From a broad range of experiments, we obtain insights into SBI’s successes and failures. Key findings include the importance of chain initialization and the increasing connectedness of modes in deeper layers. Moving beyond classical measures unfit to assess BNN inference, we develop novel convergence diagnostics that incorporate heterogeneity of layer-specific variances. As a result, our work can be seen as an extension of recent investigations into BNN symmetries (Wiese et al., 2023), prior influence (Fortuin et al., 2022), and convergence diagnostics (Vehtari et al., 2021). Our investigations reveal what BNN posteriors demand from practical SBI, leading us to propose Bayesian deep ensembles (BDEs) as a straightforward approach. BDEs achieve competitive performance and uncertainty quantification in a number of tasks.
@@ 2 Background and Related Literature
2.1 Posterior Landscape
The non-identifiability of deep neural networks (DNNs) makes inference highly challenging (Wei et al., 2023). The weight space contains multiple equioutput states, i.e., different parameter vectors leading to the same functional mapping (Hecht-Nielsen, 1990). This phenomenon generates symmetries111 We focus on what Villar et al. (2023) call passive symmetries, arising purely from modeling choices. In contrast, a separate field of research studies active symmetries, typically in the quest of making functions equivariant to symmetric properties of the physical world (e.g., Cohen & Welling, 2014). in the BNN posterior: given equal prior probability, all equioutput parameters have equal posterior density, making the posterior highly multimodal (Wiese et al., 2023). Being equioutput induces an equivalence relation (e.g., Kůrková & Kainen, 1994):
@@ 2.1 Posterior Landscape
Equioutput parameters are related by homeomorphisms $\mathcal{F}:\Theta\rightarrow\Theta$ (Grigsby et al., 2023). Depending on the network architecture, there is a huge number of (nontrivial) transformations $\mathcal{F}$ that preserve the input-output mapping, mainly arising from neuron permutability in hidden layers and certain activation functions (Chen et al., 1993; Kunin et al., 2021). Symmetries have been studied at length for DNNs optimized by empirical risk minimization (ERM). Building on pioneering work on MLPs with odd activation functions (Sussmann, 1992; Chen et al., 1993; Albertini & Sontag, 1994; Kůrková & Kainen, 1994), the rise of ReLU sparked follow-up research for homogeneous activations (e.g., Neyshabur et al., 2015; Freeman & Bruna, 2017; Bona-Pellissier et al., 2021; Grigsby et al., 2023). More recently, symmetries have been rediscovered by efforts to understand the learning dynamics of DNN optimization (Brea et al., 2019; Ainsworth et al., 2023; Entezari et al., 2022), where the closely related phenomenon of overparameterization—enabling equioutput states in the first place—plays an important role (Simsek et al., 2021; Bubeck & Sellke, 2022). Insights into the topology of DNN loss landscapes suggest the existence of multiple local optima that seem to be connected by low-loss areas for sufficiently benign curvature, depending on the amount of overparameterization and regularization present (Kuditipudi et al., 2019; Pittorino et al., 2022; Sharma et al., 2023a; Farrugia-Roberts, 2023).
Interestingly, the implications of symmetries on Bayesian inference have scarcely been discussed (with a few notable exceptions, e.g., Moore, 2016; Pourzanjani et al., 2017). Since there is a well-established link between ERM with regularized objectives and Bayesian inference222 This implies that the ERM loss is a function of the likelihood $p(\mathcal{D}|\bm{\theta})$, which is true for common loss functions like $L2$ or cross-entropy (see, e.g., Murphy, 2022). Then, we have an immediate correspondence via $p(\bm{\theta}|\mathcal{D})\propto p(\mathcal{D}|\bm{\theta})$. , the above phenomena translate from loss to posterior landscapes. In particular, we consider the following notion:
@@
Proposition 2.3 follows immediately from Bayes’ theorem and the fact that equioutput parameters share the same likelihood. Equioutput parameter states thus induce posterior multimodality, where the connectivity of those modes depends on the intricate relationship between prior belief and observed data in an overparameterized system, as discussed in the introductory example of Fig. 1. It turns out that this mode-connectedness directly affects the success of classical SBI samplers in traveling between modes.
2.2 Sampling-Based Inference
Research on SBI so far seems to disregard symmetries altogether. Hamiltonian Monte Carlo (HMC; Neal, 2011) is the de facto gold standard (Farquhar et al., 2020; Izmailov et al., 2021). Combinations of SBI with classical deep learning elements have been proposed for scalability, such as stochastic-gradient MCMC variants (Mandt et al., 2017; Zhang et al., 2020; Cobb & Jalaian, 2021), minibatch MCMC with blockwise sampling (Papamarkou, 2023), subsampling of likelihoods (Goan et al., 2023), or MCMC boosted with momentum-based optimization (Bieringer et al., 2023). Another strand of research attempts to find better priors than the popular isotropic Gaussian (Fortuin et al., 2022; Kim et al., 2023).
@@ 3 Experimental Setup
In Sections 47, we highlight findings from an extensive grid of experiments with more details and results in Appendix C. Our code is available under this URL.  Section 4 examines the general feasibility of SBI, focusing on whether SBI can achieve SOTA performance. Section 5 investigates how the multimodality of posteriors influences SBI performance, after which Section 6 discusses ways of handling multimodality to make SBI more practical. Section 7 summarizes the findings and proposes BDEs as a competitive solutions for SBI in BNNs.
Data sets and models. We consider several regression data sets from the UCI benchmark (Dua & Graff, 2017): airfoil, concrete, energy and yacht, as well as two larger data sets, bikesharing and protein. The MLPs considered are of varying width (up to 64 hidden units) and depth (up to seven layers). We investigate not only differences in model architecture, the activation function and the weight priors, but also in the choice of the sampler and its configuration.
Performance metrics. BNN predictions are computed via (empirical) Bayesian model averaging. In particular, each posterior sample induces a model from which we obtain the parameterization of a PPD conditioned on a given test observation from a test set $\mathcal{D}_{\text{test}}\in(\mathcal{X}\times\mathcal{Y})^{n_{\text{test}}}$ (20% of the data in all experiments). As a point estimate, we compute the expectation over those conditional distributions via Monte Carlo integration. Since we focus on regression, we use the test root mean squared error (RMSE) to assess predictive performance. The quality of uncertainty quantification can be measured by the log-PPD (LPPD; Eq. 7 in Appendix B) on test data333 Meaningful inference should yield (L)PPDs under which the true label $\bm{y}^{\ast}$ has high posterior density. Intuitively, this can only be the case for appropriate PPD dispersion (too concentrated—$\bm{y}^{\ast}$ will frequently not coincide with the PPD mode and thus have low density; too diffuse—no value is assigned high density). . We also assess the calibration of the samples by comparing nominal and empirical coverage of credibility intervals of the PPD.
@@
airfoil 0.716 0.288 0.252 0.239 0.250 0.182
bikesharing 0.790 0.304 0.374 0.365 0.362 0.253
concrete 0.630 0.312 0.317 0.282 0.554 0.258
energy 0.274 0.051 0.048 0.043 0.062 0.037
protein 0.863 0.700 0.804 0.803 1.077 0.716
@@ 4 General Feasibility
We hypothesize that published results obscure the fact that samplers cannot deal with all existing BNN posterior problems off-the-shelf. As a sanity check, we match the RMSE performance of SBI against that of a linear model (LM). Using the sampled chains that outperform the LM, we then compare the resulting BNN performance, i.e., the ensemble of all chains and their 8,000 samples, to a random forest (RF; with default settings) as a stronger baseline, a (non-Bayesian) DNN of the same architecture, and a deep ensemble (DE; Lakshminarayanan et al., 2017) with 12 members. The DNN and DE models are trained with Adam (Kingma & Ba, 2015, further details in Appendix D) and the BNNs are sampled using NUTS with unit Gaussian prior, and tanh activation. In order to show the ensemble effect of the BNN, we also report the performance of the model induced by a single random sample (RS) from the chains of the BNN.
General feasibility and tanh SOTA performance. Using chains that outperform the LM, samplers, in fact, yield SOTA performance across different experimental settings for tanh activation (cf. Table 1). While individual samples (RS) cannot outperform the RF, DNN or DE, their ensemble supersedes the baselines in performance. We discuss the benefits of multi-chain SBI in Sections 5 and 7.
The unboundedness problem. In contrast to the results of tanh networks, off-the-shelf SBI does not produce meaningful results for ReLU-, SiLU- or LeakyReLU-activated BNNs across different architectures, prior choices, and data sets. Instead, we observe what we call dying samplers, getting stuck almost immediately and never leaving the area of the starting value (cf. Section C in the Appendix). This becomes evident in Table 2 reporting the share of chains that perform better than an LM in terms of RMSE for ReLU and tanh activation. The number of successful chains with ReLU is notably smaller compared to tanh, which holds true also for other (unbounded) activation functions (cf. Fig. 14 in the Appendix). As is known from classical DNN optimization, weight variances in networks with unbounded activation functions can explode if not properly initialized (see, e.g., He et al., 2015). Similarly, if priors remain constant across layers with variance independent of the size of each layer, we observe that SBI will not produce chains better than a simple baseline. Hence, choosing bounded activation functions might help to achieve good performance. However, even for tanh, Table 2 shows that the more complex the architecture, the less reliable the samplers become. We elaborate on a potential solution of this problem in Section 7.
@@
Superiority of NUTS. NUTS outperforms HMC across different data sets and activation functions (cf. Table 5 in the Appendix). Overall, HMC with fixed hyperparameters can only produce chains with better performance than the LM in 1% of all experiments, suggesting that the refined handling of sampling trajectories by NUTS is indispensable for BNNs. Tuning HMC’s hyperparameters, in turn, might be less efficient than NUTS, which seems to perform well out of the box in many cases.

### v2

@@ Abstract
A major challenge in sample-based inference (SBI) for Bayesian neural networks is the size and structure of the networks’ parameter space. Our work shows that successful SBI is possible by embracing the characteristic relationship between weight and function space, uncovering a systematic link between overparameterization and the difficulty of the sampling problem. Through extensive experiments, we establish practical guidelines for sampling and convergence diagnosis. As a result, we present a deep ensemble initialized approach as an effective solution with competitive performance and uncertainty quantification.
@@ 1 Introduction
Bayesian neural networks (BNNs) represent a principled solution to the problem of probabilistic deep learning. In the absence of analytically tractable solutions, Bayesians traditionally rely on sample-based inference (SBI) as it can in theory recover the true posterior—up to a Monte Carlo error—and requires no (restrictive) assumptions on the posterior distributional family. Consequently, SBI has enormous potential for uncertainty quantification in BNNs (Farquhar et al., 2020; Izmailov et al., 2021; Wiese et al., 2023). However, recent research has focused on local approximations that may not fully capture the multimodality of BNN posteriors, as noted in Alexos et al. (2022); Arbel et al. (2023). Such approaches potentially overlook significant portions of the posterior density, precluding a comprehensive quantification of uncertainty. The main reason behind the reluctance to adopt SBI appears to be its perceived computational demands, as discussed in Papamarkou et al. (2022); Sharma et al. (2023b). Indeed, Izmailov et al. (2021) show that SBI is possible for large state-of-the-art (SOTA) networks but comes at a high cost. Yet, scalable software solutions (e.g., JAX, Bradbury et al., 2018) and methodological adaptations (Nemeth & Fearnhead, 2021) have made SBI increasingly accessible.
While technological progress continues unabated, work studying the peculiarities of SBI remains limited. One reason for this research gap, and possibly the most fundamental problem for BNNs, is their large number of parameters. The resulting challenges are uncharted territory for established Bayesian workflow routines such as the sampling strategy, the choice of prior, and convergence monitoring. It is tempting to conclude that parameter space inference should simply be discarded in favor of addressing uncertainties directly in the function space (see, e.g., Tran et al., 2022), but this perspective risks suspending valuable insights.
@@
Our Contributions. Among the discussions about the merits of parameter space inference (see also Appendix A), our paper sheds light on how SBI can be successfully leveraged for BNNs. We argue that accounting for the idiosyncratic relationship between weight and function space is key to affordable SBI. From a broad range of experiments, we obtain insights into SBI’s successes and failures. Key findings include the importance of chain initialization and the increasing connectedness of modes in deeper layers. Moving beyond classical measures unfit to assess BNN inference, we develop novel convergence diagnostics that incorporate heterogeneity of layer-specific variances. As a result, our work can be seen as an extension of recent investigations into BNN symmetries (Wiese et al., 2023), prior influence (Fortuin et al., 2022), and convergence diagnostics (Vehtari et al., 2021). Our investigations reveal what BNN posteriors demand from practical SBI, leading us to propose Deep Ensemble Initialized MCMC (DEI-MCMC) as a straightforward approach. DEI-MCMC achieves competitive performance and uncertainty quantification in a number of tasks.
@@ 2 Background and Related Literature
2.1 Multimodality in the Posterior Landscape
The non-identifiability of deep neural networks (DNNs) makes inference highly challenging (Wei et al., 2023). The weight space contains multiple equioutput states, i.e., different parameter vectors leading to the same functional mapping (Hecht-Nielsen, 1990). This phenomenon generates symmetries111 We focus on what Villar et al. (2023) call passive symmetries, arising purely from modeling choices. In contrast, a separate field of research studies active symmetries, typically in the quest of making functions equivariant to symmetric properties of the physical world (e.g., Cohen & Welling, 2014). in the BNN posterior: given equal prior probability, all equioutput parameters have equal posterior density, which leads to strong multimodality of the posterior (Wiese et al., 2023). Being equioutput induces an equivalence relation (e.g., Kůrková & Kainen, 1994):
@@ 2.1 Multimodality in the Posterior Landscape
Equioutput parameters are related by homeomorphisms $\mathcal{F}:\Theta\rightarrow\Theta$ (Grigsby et al., 2023). Depending on the network architecture, there is a huge number of (nontrivial) transformations $\mathcal{F}$ that preserve the input-output mapping, mainly arising from neuron permutability in hidden layers and certain activation functions (Chen et al., 1993; Kunin et al., 2021). Symmetries have been studied at length for DNNs optimized by empirical risk minimization (ERM). Building on pioneering work on MLPs with odd activation functions (Sussmann, 1992; Chen et al., 1993; Albertini & Sontag, 1994; Kůrková & Kainen, 1994), the rise of ReLU sparked follow-up research for homogeneous activations (e.g., Neyshabur et al., 2015; Freeman & Bruna, 2017; Bona-Pellissier et al., 2021; Grigsby et al., 2023). More recently, symmetries have been rediscovered in efforts to understand the learning dynamics of DNN optimization (Brea et al., 2019; Ainsworth et al., 2023; Entezari et al., 2022), where the closely related phenomenon of overparameterization—enabling equioutput states in the first place—plays an important role (Simsek et al., 2021; Bubeck & Sellke, 2022). Interestingly, symmetries have scarcely been discussed from a Bayesian viewpoint (with a few notable exceptions, e.g., Moore, 2016; Pourzanjani et al., 2017). Since there is a well-established link between ERM with regularized objectives and Bayesian inference, the above phenomena translate from loss to posterior landscapes (e.g., MacKay, 1992). In particular, we consider the following notion:
@@
Proposition 2.3 follows immediately from Bayes’ theorem and the fact that equioutput parameters share the same likelihood by definition.
2.2 Mode Connectivity
The existence of multiple modes in the posterior (or loss) landscape begs the question of how inference (or optimization) is affected. Loosely speaking, a multimodal surface with strong curvature makes inference harder because the sampler will take a long time to travel between modes through areas of low posterior probability, or even infinitely long in the worst case of wholly disconnected modes. The same intuition holds for standard gradient-based optimizers that risk getting stuck in isolated local optima. As a consequence, the phenomenon of mode connectivity (Garipov et al., 2018) has received increased attention in the past few years. Fortunately, local optima generally seem to be connected by low-loss areas for sufficiently benign curvature of the loss hypersurface (Kuditipudi et al., 2019; Pittorino et al., 2022; Sharma et al., 2023a; Farrugia-Roberts, 2023). Draxler et al. (2018) even conjectured that there might in fact be a single manifold containing all optima. However, it is crucial to understand that connectivity is a relative notion in three regards. First, modes will rarely be disconnected in the actual sense of the word, meaning that areas between them might be assigned very low, but not zero, posterior probability. Second, as discussed in the introductory case of Fig. 1, the curvature of the landscape is modulated by the relationship between prior belief (or inductive biases; e.g., Garipov et al., 2018) and observed data (e.g., Draxler et al., 2018). Weak priors together with a small amount of training data, for instance, will produce a rather smooth surface with less pronounced modes in absence of strong evidence for any particular hypothesis. Moreover, Ainsworth et al. (2023) suggest that mode connectivity is inextricably tied to the implicit regularization of gradient-based optimizers. Third, and perhaps most importantly for the present work, theoretical mode connectivity is not all that matters. The behavior and mobility of a sampler depend directly on its design (e.g., the propensity of accepting proposals for sampling locations). If the sampler fails to traverse low-density areas, modes become effectively disconnected even if the theory suggests the existence of a non-zero-probability link between them. In addition to the data, inductive biases and regularization, algorithmic uncertainty of the sampling process thus also plays a role.
2.3 Inference Methods
Ensemble Approaches. To account for multiple basins of attraction in the posterior landscape, several inference methods using ensembles have been proposed. Inspired by explicit ensembling of DNNs (Lakshminarayanan et al., 2017), approaches grounded more firmly in a Bayesian mindset include modifications of the forward pass to (asymptotically) obtain samples from the Gaussian process posterior (He et al., 2020) and ensembles of Gaussian posterior approximators (MultiSWAG; Wilson & Izmailov, 2020).
Sampling-Based Inference. Research on SBI, on the other hand, is largely disregarding symmetries. Hamiltonian Monte Carlo (HMC; Neal, 2011), struggling with multimodality just like any other MCMC method, is the de facto gold standard (Farquhar et al., 2020; Izmailov et al., 2021). Combinations of SBI with classical optimization elements have been proposed for scalability, such as stochastic-gradient MCMC variants (Mandt et al., 2017; Zhang et al., 2020; Cobb & Jalaian, 2021), minibatch MCMC with blockwise sampling (Papamarkou, 2023), subsampling of likelihoods (Goan et al., 2023), or MCMC boosted with momentum-based information (Bieringer et al., 2023). A separate strand of research attempts to find more informative priors (Fortuin et al., 2022; Kim et al., 2023).
@@ 3 Experimental Setup
airfoil 0.716 0.255 0.252 0.239 0.250 0.182
bikesharing 0.790 0.231 0.374 0.365 0.362 0.253
concrete 0.630 0.304 0.317 0.282 0.554 0.258
energy 0.274 0.050 0.048 0.043 0.062 0.037
protein 0.863 0.581 0.804 0.803 1.077 0.716
@@
In Sections 47, we highlight findings from an extensive grid of experiments with more details and results in Appendix C. Our code is available under this URL.  Section 4 examines the general feasibility of SBI, focusing on whether SBI can achieve SOTA performance. Section 5 investigates how the multimodality of posteriors influences SBI performance, after which Section 6 discusses ways of handling multimodality to make SBI more practical. Section 7 summarizes the findings and proposes DEI-MCMC as a competitive solutions for SBI in BNNs.
Data sets and models. We consider several regression data sets from the UCI benchmark (Dua & Graff, 2017): airfoil, concrete, energy and yacht, as well as two larger data sets, bikesharing and protein. The MLPs considered vary in width (up to 64 hidden units) and depth (up to seven layers). We investigate not only differences in model architecture, the activation function and the weight priors, but also in the choice of the sampler and its configuration.
Performance metrics. BNN predictions are computed via (empirical) Bayesian model averaging. In particular, each posterior sample induces a model from which we obtain the parameterization of a PPD conditioned on a given test observation from a test set $\mathcal{D}_{\text{test}}\in(\mathcal{X}\times\mathcal{Y})^{n_{\text{test}}}$ (20% of the data in all experiments). As a point estimate, we compute the expectation over those conditional distributions via Monte Carlo integration. Since we focus on regression, we use the test root mean squared error (RMSE) to assess predictive performance. The quality of uncertainty quantification can be measured by the log-PPD (LPPD;  Eq. 7 in Appendix B) on test data222 Meaningful inference should yield (L)PPDs under which the true label $\bm{y}^{\ast}$ has high posterior density. Intuitively, this can only be the case for appropriate PPD dispersion (too concentrated—$\bm{y}^{\ast}$ will frequently not coincide with the PPD mode and thus have low density; too diffuse—no value is assigned high density). . We also assess the calibration of the samples by comparing nominal and empirical coverage of credibility intervals of the PPD.
@@ 4 General Feasibility
We hypothesize that published results obscure the fact that samplers cannot deal with all existing BNN posterior problems off-the-shelf. As a sanity check, we match the RMSE performance of SBI against that of a linear model (LM). Using the sampled chains that outperform the LM, we then compare the resulting BNN performance, i.e., the ensemble of all chains and their 8,000 samples, to a random forest (RF; tuned with optuna (Akiba et al., 2019)) as a stronger baseline, a (non-Bayesian) DNN of the same architecture, and a deep ensemble (DE; Lakshminarayanan et al., 2017) with 12 members. The DNN and DE models are trained with Adam (Kingma & Ba, 2015, further details in Appendix D) and the BNNs are sampled using NUTS with unit Gaussian prior and tanh activation. In order to show the ensemble effect of the BNN, we also report the performance of the model induced by a single random sample (RS) from the chains of the BNN.
General feasibility and tanh SOTA performance. Using chains that outperform the LM, samplers indeed yield SOTA performance across different experimental settings for tanh activation (cf. Table 1). While individual samples (RS) cannot outperform the RF, DNN or DE, their ensemble supersedes the baselines in performance. We discuss the benefits of multi-chain SBI in Sections 5 and 7.
The unboundedness problem. In contrast to the results of tanh networks, off-the-shelf SBI does not produce meaningful results for ReLU-, SiLU- or LeakyReLU-activated BNNs across different architectures, prior choices, and data sets. Instead, we observe what we call dying samplers, getting stuck almost immediately and never leaving the area of the starting value (cf. Appendix C). This becomes evident in Table 2 reporting the share of chains that perform better than an LM in terms of RMSE for ReLU and tanh activation. The number of successful chains with ReLU is notably smaller compared to tanh, which holds true also for other (unbounded) activation functions (cf. Fig. 14 in the Appendix). As is known from classical DNN optimization, weight variances in networks with unbounded activation functions can explode if not properly initialized (see, e.g., He et al., 2015). Similarly, if priors remain constant across layers with variance independent of the size of each layer, we observe that SBI will not produce chains better than a simple baseline. Hence, choosing bounded activation functions might help to achieve good performance. However, even for tanh, Table 2 shows that the more complex the architecture, the less reliable the samplers become. We elaborate on a potential solution of this problem in Section 7.
Superiority of NUTS. NUTS outperforms HMC across different data sets and activation functions (cf. Table 5 in the Appendix). Overall, HMC with fixed hyperparameters can only produce chains with better performance than the LM in 1% of all experiments, suggesting that the refined handling of sampling trajectories by NUTS is indispensable for BNNs. Tuning HMC’s hyperparameters, in turn, might be less efficient than NUTS, with the latter performing well out of the box in many cases.